# B-cos LM: Efficiently Transforming Pre-trained Language Models for Improved Explainability

### v1

@@ 3.2.2 Fine-tuning
The B-cos adaptations above modify the architecture and computation of models, requiring fine-tuning to restore their capabilities and adapt to alignment pressure. Following the NLP-typical “pre-train then fine-tune” paradigm, we directly transform PLMs to B-cos LMs, rather than adapting task-specific models as done in previous work (Arya et al., 2024). This fundamental difference in the training pipeline adds complexity to B-cosification in NLP, as the objective involves both B-cos conversion and task fine-tuning. While there are multiple ways to conjoin these two steps (cf. § 7), we find that the most efficient way is to combine them by first applying B-cos adaptations to a PLM and then fine-tuning it on a downstream task. Following Böhle et al. (2022), we use the binary cross-entropy (BCE) loss instead of conventional cross-entropy loss, as it explicitly maximizes the absolute target logits and strengthens alignment pressure. We provide an extensive comparison of B-cosification setups in § 7.
@@ 3.3 Computing B-cos Explanations
Once trained, the B-cos LM can generate explanations that faithfully summarize its decision-making process during inference. As all components are dynamic linear with no bias terms (cf. Appendix D), the entire model computation can be expressed as a sequence of dynamic linear transformations:
$\mathbf{\hat{W}}_{L}(\mathbf{A}_{L})\mathbf{\hat{W}}_{L-1}(\mathbf{A}_{L-1})...\mathbf{\hat{W}}_{1}(\mathbf{A}_{1}=\mathbf{X})\mathbf{X}$ (3)
which can be completely summarized as a single dynamic linear function $\Pi_{j=1}^{L}\mathbf{\hat{W}}_{j}(\mathbf{A}_{j})$. [^3] Note that a residual connection of $\mathbf{W(x)x+x}$ with $\mathbf{x}\in\mathbb{R}^{n}$ and $\mathbf{W(x)}\in\mathbb{R}^{n\times n}$ is mathematically equivalent to a single dynamic linear transformation of $(\mathbf{W(x)+I}_{n})\mathbf{x}$. Considering the textual inputs specific to NLP, we attribute the model’s predictions to the embedding representations. Specifically, to quantify the contribution of a token $i$ to a model prediction, we compute the dot product $\mathbf{W(x}_{i})\mathbf{x}_{i}$ between its embedding $\mathbf{x}_{i}$ and the corresponding dynamic linear weight $\mathbf{W(x}_{i})$ for the predicted class logit. For the remainder of the paper, we will refer to such explanations as B-cos explanations.
Model Method AG News IMDB HateXplain
(a) Conv. BERT Attention 24.40 8.09 50 26.84 14.56 50 27.64 13.83 50
IxG 15.28 10.19 45.41 18.29 16.96 49.42 19.16 18.90 47.24
SIG 27.02 3.40 64.77 29.34 14.05 59.09 37.31 5.10 66.38
DecompX 52.16 0.92 84.48 57.94 2.41 63.27 44.86 2.72 66.76
ShapSampl 43.96 0.46 82.87 58.29 2.44 71.29 44.86 2.43 67.17
LIME 44.95 0.06 80.28 51.45 6.07 60.15 22.64 14.30 57.61
(b) B-cos BERT Attention 33.51 2.71 50 32.91 5.31 50 39.3 3.93 50
IxG 26.89 1.31 48.24 61.74 -2.39 52.95 44.93 -0.60 53.57
SIG 14.39 4.65 19.64 29.06 3.11 56.82 35.04 1.96 60.23
DecompX - - - - - - - - -
ShapSampl 15.90 3.91 52.71 35.95 0.66 52.14 39.6 0.66 65.02
LIME 57.99 0.07 79.30 70.05 -1.55 60.03 40.84 3.84 59.14
(c) B-cos BERT B-cos 64.22 -1.26 87.92 75.33 -2.95 70.27 59.66 -4.89 77.57
Table 2: Faithfulness evaluation for conventionally fine-tuned BERT and B-cos BERT across three datasets. The best results are in bold. We find that B-cos explanations are consistently more faithful than post-hoc explanations from both models.
@@ 4 Experiments
We evaluate the task performance of B-cos LMs and faithfulness of B-cos explanations against conventional models and baseline explanation methods across various tasks, PLMs, and metrics.
@@
Our experiments include three sequence classification datasets: AG News (topic classification, Zhang et al., 2015), IMDB (sentiment analysis, Maas et al., 2011), and HateXplain (hate speech detection, Mathew et al., 2021). We use BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), and DistilBERT (Sanh et al., 2019) as the basis for conventional fine-tuning and for obtaining B-cos LMs (B=1.5) with the same training hyperparameters (cf. Appendix E for details on fine-tuning, B-cosification, and data splits).
@@
Compared to perturbation-based metrics, SeqPG does not rely on perturbations and thus avoids the potential distortions introduced by token masking. When constructing SeqPG examples, we truncate each segment to a fixed length and randomize segment order to control for length and position effects. We generate synthetic examples using correctly and most confidently classified test instances. SeqPG can be seen as a standardized version of hybrid document evaluation (Poerner et al., 2018). We provide an example of SeqPG in Figure 7 and more details in Appendix F.
Baselines
We compare B-cos explanations against a diverse set of post-hoc explanation methods: Attention (Bahdanau et al., 2015), InputXGradient (IxG, Kindermans et al., 2016), Sequential Integrated Gradients (SIG, Enguehard, 2023), DecompX (Modarressi et al., 2023), Shapley Value Sampling (ShapSampl, Strumbelj and Kononenko, 2010), and LIME (Ribeiro et al., 2016). For a fair comparison against embedding-level methods, we aggregate attributions by summing across all embedding dimensions (cf. Appendix E).
Figure 2: Mean accuracy of conventionally fine-tuned and B-cos BERT models averaged over three runs. B-cos models perform comparably to conventional models on most tasks.
@@
Figure 2 shows the accuracy of conventionally fine-tuned and B-cos BERT across three datasets (we provide results for DistilBERT and RoBERTa in Appendix G). We find that all B-cos LMs performs on par with conventional models on AG News and HateXplain, with only a minor drop ($\sim$1%) in accuracy. Only for IMDB, we find a slightly larger drop of 4.21%, though the performance remains strong overall.
@@
Table 2 shows the faithfulness scores for post-hoc explanation methods on (a) conventionally fine-tuned BERT models and (b) B-cos BERT models, as well as (c) B-cos explanations extracted from B-cos BERT. The results show that B-cos explanations are consistently and substantially more faithful than post-hoc methods across all datasets. This improvement holds both across different models and within the same model. B-cos explanations outperform the strongest post-hoc methods on conventional models by an average of 14.63 points in Comp score and achieve negative Suff scores, indicating that the identified important tokens alone enable even more confident predictions. Additionally, B-cos explanations show a considerable improvement in SeqPG. Similar trends are observed for DistilBERT and RoBERTa (Appendix H), further strengthening our findings.
@@
Figure 3: Human evaluation reveals that B-cos explanations have better human interpretability and human agreement than baseline methods and the improvements are statistically significant.
5 Human Evaluation
We conduct a human study to evaluate the human interpretability and agreement of B-cos explanations, comparing them against three strong post-hoc explanation methods on the conventional BERT model. For the study, we randomly select 50 instances from AG News and HateXplain where the B-cos and conventional model predict the same label. We then ask five annotators to rate the respective explanations in terms of human interpretability (how well they understand it) and human agreement (how much they agree with the it) on a scale of 1-5. We provide further details of the human evaluation in Appendix I.
Human Evaluation Results
@@ 5 Human Evaluation
Figure 3 shows that B-cos explanations have a better human interpretability and exhibit greater alignment with human reasoning than post-hoc methods. Conducting paired t-tests with a Bonferroni-corrected $\alpha=\frac{0.05}{6}=0.008\overline{3}$ (Bonferroni, 1936) shows that the improvements of B-cos explanations are statistically significant ($p<\alpha$) for both metrics.
6 Qualitative Analysis
Figure 4: Examples of B-cos explanations (B-cos BERT) as well as ShapSampl and DecompX explanations (BERT) from AG News. Green (red) indicates the positive (negative) impact of tokens on the prediction. The B-cos explanation highlights only relevant tokens and is more interpretable to humans (cf. Appendix J for more examples).
@@ 6 Qualitative Analysis
Figure 4 provides an example of B-cos and other (post-hoc) explanations. It can be seen the B-cos explanation highlights important tokens well with little focus on irrelevant ones. In contrast, ShapSampl attributes the highest importance to the [SEP] token and provides only little useful information. Meanwhile, DecompX extracts a significant amount of irrelevant information. Overall, we find that the B-cos explanation provides clearer and more relevant attributions compared to the post-hoc explanations.
@@ 7 Comparison of B-cosification Setups
  • Task then B-cos: PLMs are first fine-tuned on a downstream task. B-cos adaptations are then applied, followed by further fine-tuning on the same task for B-cos conversion. This setup is equivalent to Arya et al. (2024) who apply B-cosification to models with task capabilities.
  • B-cos then task: B-cos adaptations are applied to PLMs first, followed by pre-training on unsupervised texts to enhance B-cosification (cf. Appendix E). The pre-trained B-cos models are then fine-tuned on the downstream task.
We evaluate these setups against the B-cosification approach used in our main experiments (B-cos LM) and compare task performance, faithfulness, and training efficiency. Additionally, we report results for conventional fine-tuning (Conv. LM) and training a randomly initialized B-cos LM (B-cos from scratch). Experiments are conducted on IMDB, with results averaged over three runs.
@@
Conv. LM 5 94.06 - 7.33
B-cos LM 5 89.85 70.41 3.67
B-cos from scratch 5 88.36 60.92 4.33
Task then B-cos 1+4 90.14 70.28 1+4.33
2+3 90.33 70.36 3+3.33
3+2 90.07 69.94 4+3
4+1 88.19 70.36 5+1
5+5* 90.33 69.65 6.67+3.33
B-cos then task 1+4 89.78 65.58 1+5.67
2+3 89.81 66.01 3+4
3+2 89.38 66.95 4+3
4+1 87.42 67.9 6+1
5+5* 90.38 71.16 7+3
10+5* 91.08 75.06 15+3.67
20+5* 91.75 76.66 31+6.33
Table 3: Different B-cosification setups. For two-phase methods, we report epoch distribution and convergence steps per phase. * marks additional training epochs.
Table 3 shows that B-cos LM requires fewer training steps to reach optimal validation performance than conventional fine-tuning. Training B-cos LM from scratch results in worse task performance and faithfulness, emphasizing the importance of good parameter initialization. Among the two setups that separate task fine-tuning and B-cos conversion, Task then B-cos achieves results comparable to B-cos LM but requires more training steps. B-cos then task initially performs worse under the same training budget. However, with additional pre-training epochs, it surpasses other B-cosification setups in both task performance and faithfulness. Overall, we find that combining task fine-tuning and B-cos conversion is the most efficient approach. However, with sufficient pre-training, B-cos then task can produce more performant and explainable models.
8 Effects of B-cosification and B Values
For a deeper understanding of how B-cosification and parameter B affect model performance and behavior, we compare conventional and B-cos BERT trained on HateXplain across different B values. We also provide an empirical analysis of the impact of B on input-weight alignment in Appendix K.

### v2

@@ 3.3.2 Fine-tuning
The B-cos adaptations above modify the architecture and computation of models, requiring fine-tuning to restore their capabilities and adapt to alignment pressure. Following the “pre-train then fine-tune” paradigm, which is frequently utilized in NLP tasks, we directly transform PLMs to B-cos LMs, rather than adapting task-specific models as done in previous work (Arya et al., 2024). This fundamental difference in the training pipeline adds complexity to B-cosification for LMs, as the objective involves both B-cos conversion and task fine-tuning. While there are multiple ways to conjoin these two steps (cf. § 5), we find that the most efficient way is to combine them by first applying B-cos adaptations to a PLM and then fine-tuning it on a downstream task. Following Böhle et al. (2022), we use the binary cross-entropy (BCE) loss instead of the conventional cross-entropy loss, as it explicitly maximizes the absolute target logits and strengthens the alignment pressure. We provide an extensive comparison of different B-cosification setups in § 5.
@@ 3.4 Computing B-cos Explanations
Once trained, the B-cos LM can generate explanations that faithfully summarize its decision-making process during inference. As all components are dynamic linear with no bias terms (cf. § 3.2), the entire model computation can be expressed as a sequence of matrix multiplications, which can be completely summarized as a single dynamic linear function:
$\mathbf{\hat{W}}_{L}(\mathbf{A}_{L})\mathbf{\hat{W}}_{L-1}(\mathbf{A}_{L-1})...\mathbf{\hat{W}}_{1}(\mathbf{A}_{1}=\mathbf{X})\mathbf{X}=\Pi_{j=1}^{L}\mathbf{\hat{W}}_{j}(\mathbf{A}_{j})$ (6)
Note that a residual connection of $\mathbf{W(x)x+x}$ with $\mathbf{x}\in\mathbb{R}^{n}$ and $\mathbf{W(x)}\in\mathbb{R}^{n\times n}$ is mathematically equivalent to a single dynamic linear transformation of $(\mathbf{W(x)+I}_{n})\mathbf{x}$. Considering the textual inputs specific to LMs, we attribute the model’s predictions to the embedding representations. Specifically, to quantify the contribution of a token $i$ to a model prediction, we compute the dot product $\mathbf{W(x}_{i})\mathbf{x}_{i}$ between its embedding $\mathbf{x}_{i}$ and the corresponding dynamic linear weight $\mathbf{W(x}_{i})$ for the target class logit. For the remainder of the paper, we will refer to such explanations as B-cos explanations.
@@ 4 Experiments
We evaluate the task performance of B-cos LMs and faithfulness of B-cos explanations with automatic evaluation across various tasks and PLMs. In addition, we conduct a human evaluation study to compare the human interpretability of B-cos explanations. §4.14.3 describe our automatic evaluation setup, results, as well as human evaluation study, respectively. §4.4 provides a qualitative analysis. Finally, we study the computational efficiency in §4.5 and conduct an ablation study in §4.6. More details on the experimental setup and baseline methods are provided in Appendix C.
4.1 Experimental Setup
@@
Our experiments use three datasets: AG News (topic classification, Zhang et al., 2015), IMDB (sentiment analysis, Maas et al., 2011), and HateXplain (hate speech detection, Mathew et al., 2021). BERT (Devlin et al., 2019), DistilBERT (Sanh et al., 2019), and RoBERTa (Liu et al., 2019) are used as the basis for conventional fine-tuning and for obtaining B-cos LMs. We set B=1.25 for IMDB and B=1.5 for AG News and HateXplain datasets.
Baselines
We compare B-cos explanations against a diverse set of post-hoc explanation methods: Attention (Bahdanau et al., 2015), InputXGradient (IxG, Kindermans et al., 2016), Sequential Integrated Gradients (SIG, Enguehard, 2023), DecompX (Modarressi et al., 2023), Shapley Value Sampling (ShapSampl, Strumbelj & Kononenko, 2010), and LIME (Ribeiro et al., 2016). We also apply these methods to a model trained with Saloss (Chrysostomou & Aletras, 2021), which incorporates additional faithfulness regularization. This setup enables a direct comparison between B-cos LMs and models specifically optimized for explainability. For embedding-level explanation methods, we aggregate attributions by summing across all embedding dimensions.
@@
Compared to perturbation-based metrics, SeqPG does not rely on perturbations and thus avoids the potential distortions introduced by token masking. When constructing SeqPG examples, we truncate each segment to a fixed length and randomize segment order to control for length and position effects. We generate synthetic examples using correctly and most confidently classified test instances. SeqPG can be seen as a standardized version of hybrid document evaluation (Poerner et al., 2018). We provide an example of SeqPG in Figure 7 and more details in Appendix D.
4.2 Automatic Evaluation Results
@@
Figure 2: Mean accuracy of conventionally fine-tuned, Saloss and B-cos BERT models averaged over three runs. We use B=1.5, 1.25, and 1.5 for AG News, IMDB, and HateXplain, respectively. B-cos models perform comparably to conventional models on most tasks.
Figure 2 shows the accuracy of conventionally fine-tuned, Saloss and B-cos BERT across three datasets (we provide results for DistilBERT and RoBERTa in Appendix E). On AG News and HateXplain, B-cos LMs performs on par with conventional models, with only a minor drop ($\sim$1%) in accuracy. They also outperform Saloss models on these datasets. Only for IMDB, we find a slightly larger drop of 3.06% compared to conventional BERT, though the performance remains strong overall.
@@
Table 2 shows the faithfulness scores for post-hoc explanation methods on conventionally fine-tuned and Saloss BERT models, as well as B-cos explanations from B-cos BERT. The results show that B-cos explanations are consistently and substantially more faithful than post-hoc methods across all models and datasets. On average, B-cos explanations outperform the strongest post-hoc methods on conventional models by 14.63 points in Comp and achieve negative Suff scores, indicating that the identified important tokens alone enable even more confident predictions. B-cos also shows significant gains in SeqPG. While Saloss improves faithfulness for some post-hoc methods over conventional models, it still underperforms compared to B-cos LMs by a large margin. Similar trends are observed for DistilBERT and RoBERTa (Appendix F) as well, further strengthening our findings. Although we do not include rationale-based models in the main experiments because they typically require additional supervision, a supplementary comparison in Appendix G shows that B-cos BERT still outperforms a rationale-based model on HateXplain.
Model Method AG News IMDB HateXplain
Conv. BERT Attention 24.40 8.09 50 26.84 14.56 50 27.64 13.83 50
IxG 15.28 10.19 45.41 18.29 16.96 49.42 19.16 18.90 47.24
SIG 27.02 3.40 64.77 29.34 14.05 59.09 37.31 5.10 66.38
DecompX 52.16 0.92 84.48 57.94 2.41 63.27 44.86 2.72 66.76
ShapSampl 43.96 0.46 82.87 58.29 2.44 71.29 44.86 2.43 67.17
LIME 44.95 0.06 80.28 51.45 6.07 60.15 22.64 14.30 57.61
Saloss BERT Attention 34.73 3.65 50 27.59 13.64 50 34.95 26.26 50
IxG 14.98 12.66 51.01 24.19 16.30 49.02 26.61 30.94 50.74
SIG 16.70 8.22 63.74 45.44 8.48 54.96 44.53 21.50 54.70
DecompX 59.37 0.30 75.34 59.42 5.38 62.02 58.71 13.23 65.17
ShapSampl 37.73 0.77 73.96 65.38 3.17 70.23 57.05 15.10 72.36
LIME 53.18 2.37 76.16 53.31 6.32 58.65 21.73 21.96 55.71
B-cos BERT B-cos 64.22 -1.26 87.92 74.18 -2.87 70.43 59.66 -4.89 77.57
Table 2: Faithfulness evaluation for conventionally fine-tuned BERT, Saloss BERT and B-cos BERT across three datasets. We use B=1.5, 1.25, and 1.5 for AG News, IMDB and HateXplain, respectively. The best results are in bold. We find that B-cos explanations are consistently more faithful than post-hoc explanations from both models.
4.3 Human Evaluation
Contrary to previous B-cos studies that rely solely on automatic evaluations to assess explanations, we conduct the first human study to better evaluate the human interpretability and agreement of B-cos explanations. We compare B-cos explanations against three strong post-hoc explanation methods on the conventional BERT model.
Following the practice in Enguehard (2023) and Yue et al. (2022), we randomly select 50 instances, respectively, from AG News and HateXplain where the B-cos and conventional models make the same prediction. Five annotators then rate the explanations in terms of human interpretability (how well they understand them) and human agreement (how much they agree with them) on a scale of 1-5. Before the annotation began, the annotators were provided with a clear description of the evaluation task, metrics, and rating scales. They were also shown example annotations along with the reasoning behind each rating to help them better understand the evaluation criteria. Further details on the evaluation criteria, rating scales, annotator instructions, and example annotations can be found in Appendix H.
@@ 4.3 Human Evaluation
Figure 3: Human evaluation reveals that B-cos explanations have better human interpretability and human agreement than baseline methods.
@@
Figure 3 shows that B-cos explanations have a better human interpretability and exhibit greater alignment with human reasoning than post-hoc methods, even though they are not directly optimized for human agreement. Paired t-tests with a Bonferroni-corrected significance level $\alpha=\frac{0.05}{6}=0.008\overline{3}$ (Bonferroni, 1936) shows that the improvements of B-cos explanations are statistically significant ($p<\alpha$) for both metrics.
4.4 Qualitative Analysis
@@
Figure 4 provides an example of B-cos and other (post-hoc) explanations. It can be seen that the B-cos explanation highlights important tokens correctly with little focus on irrelevant ones. In contrast, ShapSampl attributes the highest importance to the [SEP] token and provides only little useful information. Meanwhile, DecompX extracts a significant amount of irrelevant information. Overall, we find that the B-cos explanation is more interpretable to humans by providing clearer and more relevant attributions compared to the post-hoc explanations.
Figure 4: Examples of B-cos explanations (B-cos BERT) as well as ShapSampl and DecompX explanations (conv. BERT) from AG News. Green (red) indicates the positive (negative) impact of tokens on the prediction. The B-cos explanation highlights only relevant tokens and is more interpretable to humans (cf. Appendix I for more examples).
4.5 Explanation Efficiency
Beyond improved faithfulness and human interpretability, B-cos explanations are also efficient to extract. Comparing their computational costs with strong post-hoc methods shows that B-cos explanations are the most efficient in both time and memory usage (Table 3). Post-hoc and B-cos explanations are generated from the conventionally fine-tuned and B-cos BERT models on IMDB, respectively.
4.6 Ablation Study
Method Time (s) Memory (GB)
ShapSampl 37.22 21.95
LIME 6.82 21.96
SIG 67.46 29.09
DecompX 0.76 48.38
B-cos 0.08 2.82
Table 3: Computational costs per example of generating explanations for 100 instances using an NVIDIA H100 GPU (batch size 1). B-cos explanations (bold) are at least 9x faster and require at most $\frac{1}{8}$ of VRAM.
To better understand B-cos LMs, we conduct an ablation study evaluating the impact of key design choices on task performance and explanation faithfulness.
In Table 4, we find that removing alignment pressure (using B=1) degrades both task performance and explanation faithfulness. Replacing cross-entropy with BCE loss has little effect on classification accuracy, but improves faithfulness in perturbation-based evaluations. Architectural adaptations, including removing bias terms and eliminating activation functions in prediction heads, are also critical for enhancing task performance and explainability. Besides, we observe numerical instability when generating explanations without these architectural adaptations, as the dynamic linear weights for sigmoid and tanh ($\text{sigmoid}(\mathbf{x})\times\mathbf{x^{-1}}$ and $\text{tanh}(\mathbf{x})\times\mathbf{x^{-1}}$) become unstable when $\mathbf{x}$ is close to zero.
In addition to ablations of model design and training components, we also evaluate alternative explanation methods. Replacing the dynamic linear weights $\mathbf{W(x)}$ with gradients (equivalent to IxG) yields less faithful explanations on B-cos LMs. Besides, directly extracting B-cos-like explanations, $\mathbf{W(x)x}$, from a conventional model results in worse faithfulness compared to those from B-cos LMs. [^3] Extracting $\mathbf{W(x)x}$ from conventional models follows the same approach as in B-cos LMs (cf. § 3.2), except that in standard linear transformations, the dynamic linear weight is replaced by the fixed weight matrix $\mathbf{W}$.
Acc ($\uparrow$) Comp ($\uparrow$) Suff ($\downarrow$) SeqPG ($\uparrow$)
Full system 78.64 59.66 -4.89 77.57
w/o alignment pressure (B=1) 78.07 (0.57) 57.19 (2.44) -2.57 (2.32) 70.18 (7.39)
w/o BCE training 79.00 (0.36) 49.22 (10.44) -7.91 (3.02) 79.21 (1.64)
w/o architectural adaptations 77.65 (0.99) 52.23 (7.43) -3.80 (1.09) 74.30 (3.27)
w/o dynamic linear weights (IxG) 78.64 (0.00) 44.93 (14.73) -0.60 (4.29) 53.57 (24.00)
$\mathbf{W(x)x}$ from conv. model 80.77 (2.13) 44.92 (14.74) 2.80 (7.69) 70.20 (7.37)
Table 4: Ablation study of key designs in the B-cos BERT model on HateXplain. Values in parentheses indicate the difference from the full model’s performance. Green (red) indicates the results are better (worse) than the full system.
@@ 5 Comparison of B-cosification Setups
  • Task then B-cos: PLMs are first fine-tuned on a downstream task. B-cos adaptations are then applied, followed by further fine-tuning on the same task for B-cos conversion. This setup is equivalent to Arya et al. (2024) who apply B-cosification to models with downstream task capabilities.
  • B-cos then task: B-cos adaptations are applied to PLMs first, followed by pre-training on unsupervised texts to enhance B-cosification. The pre-trained B-cos models are then fine-tuned on the downstream task.
We evaluate these setups against the B-cosification approach used in our main experiments (B-cos LM) and compare task performance, faithfulness, and training efficiency (cf. Appendix C for B-cos pre-training details). Additionally, we report results for conventional fine-tuning (Conv. LM) and training a randomly initialized B-cos LM (B-cos from scratch). Experiments are conducted on IMDB with B=1.25 for B-cos models, with results averaged over three runs.
@@
Conv. LM 5 94.06 - 6.67
B-cos LM 5 91.00 70.66 4.33
B-cos from scratch 5 88.25 60.92 4.33
Task then B-cos 1+4 91.17 70.01 1+5
2+3 91.30 70.48 3+3.33
3+2 91.38 70.83 4+3
4+1 89.56 70.66 5+1
5+5* 91.27 70.78 6.67+3.33
B-cos then task 1+4 90.64 67.07 1+5
2+3 91.04 68.97 3+4
3+2 90.50 68.48 4+3
4+1 89.18 69.92 6+1
5+5* 91.45 71.86 7+5.33
10+5* 92.19 73.44 15+6.33
20+5* 92.87 75.01 31+6
Table 5: Training epochs, accuracy, explanation faithfulness, and convergence steps for different B-cosification setups. For two-phase methods, we report epoch distribution and convergence steps per phase. * marks additional training epochs.
Table 5 shows that B-cos LM requires fewer training steps to reach optimal validation performance than conventional fine-tuning. Training B-cos LM from scratch results in worse task performance and faithfulness, emphasizing the importance of good parameter initialization. Among the two setups that separate task fine-tuning and B-cos conversion, Task then B-cos achieves results similar to B-cos LM but requires more total training steps. B-cos then task initially performs worse under the same training budget. However, with additional pre-training epochs, it surpasses other B-cosification setups in both task performance and faithfulness. Overall, we find that combining task fine-tuning and B-cos conversion is the most efficient approach. However, with sufficient pre-training, B-cos then task can produce more performant and explainable models.
6 Impact of B-cosification and B Values
For a deeper understanding of how B-cosification and alignment pressure parameter B affect model performance and behavior, we compare conventional and B-cos BERT trained on HateXplain across different B values. We also provide an empirical analysis of the impact of B on input-weight alignment in Appendix J.